# Engineering Carbon Credits Towards A Responsible FinTech Era: The Practices, Implications, and Future

### v1

@@ 5 Algorithms for Predicting Carbon Credit Prices
[52] LS-SVM Improved Particle Swarm Optimization (IPSO) EMD Y
[54] LSTM, GRU, MLP, BPNN Genetic Algorithm (GA) CEEMDAN+Permutation Entropy (PE)+Variational Mode Decomposition (VMD) N
[56] KELM Multi-Objective Chaotic Sine Cosine Algorithm (MOCSCA) ICEEMDAN Y
[59] SVR (high-frequency subsequences) + MLP (low-frequency subsequences) - ICEEMDAN+Discrete Wavelet Transform (DWT) Y
[62] BIGRU point prediction + GPR interval prediction - CEEMDAN+Wavelet Transform (WT) Y
[65] GRU - Empirical Wavelet Transform (EWT) N
[67] Extreme Learning Machine, ELM Cooperative Kidney Algorithm (CKA) Mean value Optimization Empirical Mode Decomposition, MOEMD Y
[70] LSTM - MEEMD Y
[76] XGBoost, RF, KNEA, SVM, RBFNN GWO, PSO, SA, FFOA CEEMDAN N
[79] Weighted Regularized Extreme Learning Machine, WRELM Multi-Objective Grasshopper Optimization Algorithm, MOGOA ICEEMDAN Y
[82] LSTM IWOA MEEMD N
[84] VGM (long-term trend), ELM (short-term trend) PSO Hodrick-Prescott, HP filter Y
[86] LSTM, Lower Upper Bound Estimation (LUBE) Asymmetric Multi-Objective Evolutionary Algorithm, MOEA CEEMDAN Y
[90] Light Gradient Boosting Machine (LightGBM) Bayesian Optimization Algorithm, BOA CEEMDAN, Two-Stage Feature Selection, TFS Y
[93] LSTM - Convolutional Neural Network (CNN) N
[95] LSTM - Multivariate Variational Mode Decomposition, MVMD Y
[99] LSTM, GRU Cuckoo Search Algorithm, CS CEEMDAN, Fuzzy Entropy, FE, Variational Mode Decomposition (VMD) Y
[101] Deep Augmented Frequency Enhanced Decomposed Transformer (DA-FEDformer), Multimodel Optimization Piecewise Error Correction System - ICEEMDAN N
Summarizing the research progress in carbon credit price prediction over the past few years, we can observe a gradual evolution and enhancement of the techniques employed. Early studies primarily relied on single models, such as the Hybrid Neural Fuzzy Controller (PATSOS), Adaptive Neuro-Fuzzy Inference System (ANFIS), and Artificial Neural Networks (ANN). These models played a crucial role in the initial stages of carbon price prediction [102]. However, as data complexity increased, researchers began introducing data decomposition techniques, such as Empirical Mode Decomposition (EMD) and its improved versions. These techniques allow time series data to be decomposed into more regular components, making it easier for machine learning algorithms to capture underlying patterns in the data. Additionally, optimization algorithms have been widely applied to enhance the overall performance of the models.
@@
To further improve prediction accuracy, recent research has concentrated on more refined data decomposition and optimized multi-model combinations. On the one hand, after initially decomposing the data, researchers have employed more precise secondary decomposition techniques to extract potential patterns, enabling the models to learn and predict more accurately. On the other hand, the combination of different prediction models and optimization algorithms has been extensively explored to identify the most suitable hybrid models for different carbon markets. Moreover, some studies have decomposed raw data into components with different frequencies and applied different prediction algorithms to each frequency component, ultimately integrating the predictions of each sub-model through weighted ensemble methods to enhance overall prediction performance.
In addition to point predictions, an increasing number of studies have also introduced interval prediction techniques to provide a prediction range for carbon prices. This approach not only improves the reliability of predictions but also better addresses the uncertainties present in the market, offering more comprehensive predictive insights.
@@ 6 Prediction of Enterprise Carbon Emissions
Corporate carbon emission prediction algorithms hold significant importance in the current global context of combating climate change. Beyond assisting organisations in better managing their carbon emissions, these predictive methods are also crucial in meeting the needs of investors and governments. Financial institutions are increasingly focused on whether organisations adhere to ESG principles and prefer to invest in those that can create long-term social value. Accurate carbon emission predictions help enhance corporate transparency, reduce information asymmetry, and strengthen market trust. Moreover, as governments impose stricter requirements on corporate carbon emission disclosures, precise carbon emission predictions can assist organisations in better complying with regulations, supporting government efforts to achieve emission reduction targets, and effectively evaluating whether organisations are truthfully disclosing and adhering to carbon emission rules. The following section will review various carbon emission prediction methods.
6.1 Empirical model:
@@
In the context of predicting corporate-level carbon emissions, the study by Goldhammer et al. presents an innovative approach that estimates corporate carbon footprints (CCFs) using externally available data. They employ a regression analysis model that primarily utilizes five key variables as input data: firm size (measured by turnover or number of employees), vertical integration, capital intensity, production centrality, and the carbon intensity of the national energy mix. The model was trained on actual CCF data from 93 European organisations across the chemical, construction engineering, and industrial machinery sectors. The study demonstrates that this method can accurately estimate corporate carbon emissions without relying on internal data, making it particularly suitable for scenarios requiring rapid, external assessments [103].
@@
Shakdwipee and Lee explores the application of two primary technical methods. The first is based on the Economic Input-Output Life Cycle Assessment (EIO-LCA) model, which utilizes macroeconomic data, such as industry GDP and carbon emission intensity, to estimate corporate carbon emissions. The second method is the estimation model developed by MSCI ESG Research, which combines historical company disclosures with industry-specific intensity to predict carbon emissions for years without disclosed data. The research used the 2013 carbon emission data from 277 organisations in the MSCI ACWI IMI index as training data and evaluated the models’ prediction accuracy by comparing them with the organisations’ actual disclosed data. The results indicate that the MSCI ESG Research model outperforms the EIO-LCA model in estimation accuracy, particularly in handling company-specific carbon emission intensities [104].
@@
The study by Griffin et al. provides an effective technical approach for estimating corporate greenhouse gas (GHG) emissions using external data. This research employs the Ohlson valuation model [105] and combines it with reported GHG emission data to develop a GHG emission estimation model through regression analysis. This model utilizes a range of company-specific variables as input data, including revenue, capital expenditures, and asset composition, to predict the emissions of organisations that have not disclosed their GHG data [106].
6.2 Machine learning method:
@@
Zhou et al. developed a carbon risk prediction model based on Support Vector Machines (SVM) to accurately predict carbon emissions for heavily polluting industrial enterprises in China. The study first established a carbon risk assessment system comprising 45 initial indicators. After performing principal component analysis and independent analysis, 18 core indicators were selected. The SVM algorithm was used for model training, and the penalty parameter C and Gaussian kernel parameter gamma were optimized through grid search and cross-validation to handle the nonlinear and high-dimensional characteristics of the data [107].
@@
Quyen et al. developed a novel approach for predicting corporate-level carbon emissions, specifically targeting organisations that do not publicly disclose their emissions data. The framework proposed a Meta-Elastic Net learning model to aggregate predictions from multiple base models, including Ordinary Least Squares (OLS), Elastic Net, Neural Networks, K-Nearest Neighbors (KNN), Random Forest (RF), and Extreme Gradient Boosting (XGBoost). The input training data comprised global carbon emissions data from 2,289 organisations between 2005 and 2017, totaling 13,435 annual observations. The data included Scope 1, Scope 2, and Scope 3 emissions, as well as various corporate factors such as operational scale, business models, technological advancements, and energy consumption. The model’s performance was validated through extensive cross-validation and robustness tests, demonstrating significant improvements in prediction accuracy compared to traditional regression models [108].
@@
Li et al. proposed an advanced carbon emission prediction model tailored for enterprise-level applications, integrating multi-dimensional feature extraction with an optimized Radial Basis Function (RBF) neural network. The model utilizes Principal Component Analysis (PCA) [109] combined with kernel functions to extract essential features from a diverse set of input data, which includes production metrics, fuel consumption, electricity and heat energy purchases, and other relevant economic indicators. This approach effectively reduces data complexity while preserving key predictive information. The extracted features are then input into an RBF neural network, whose parameters are finely tuned using a hybrid optimization method that combines the Drosophila Optimization Algorithm (DOA) and Differential Evolution (DE). This dual optimization strategy significantly enhances the accuracy and robustness of the emission predictions [110].
@@
Chen et al. proposed an efficient Fast Transfer-Based Adaptive Recurrent Neural Network (SEER) model aimed at addressing the high computational costs associated with real-time predictions in existing deep learning models. The SEER model enhances prediction accuracy and training efficiency by integrating data preprocessing, feature extraction, attention mechanisms, transfer learning, and adaptive learning rate mechanisms. Using carbon emission and air quality data from Shanghai and Beijing as inputs, the model effectively identifies key data points and optimizes generalization capabilities, demonstrating its accuracy and applicability in corporate carbon emission predictions [111].
@@
Tang et al. proposed CarbonNet, a novel approach for enterprise-level carbon emission prediction, aimed at overcoming the limitations of traditional macro-level forecasting models. The paper first constructs a large-scale dataset encompassing 3346 organisations over 31 years, integrating carbon emission data with corporate financial and operational metrics. To enhance prediction accuracy, the authors employed Factor Analysis for feature extraction, reducing the impact of irrelevant variables, thereby building a more generalizable carbon emission prediction model. Subsequently, the paper utilized the XGBoost algorithm for large-scale data mining, iteratively optimizing prediction errors, resulting in a median absolute relative error (MARE) of 0.25, which represents a 22% improvement in prediction accuracy compared to existing methods [112].
@@
Shou developed a corporate carbon performance evaluation framework using Support Vector Machine (SVM) to predict carbon emissions at the enterprise level. The framework, named Enterprise Carbon Performance Index System (ECPIS), integrates SVM with real-time data collection and monitoring to assess and forecast carbon emissions, enabling organisations to make data-driven decisions to reduce their carbon footprint. The input training data for the SVM model includes historical carbon emissions data, energy consumption metrics, operational data such as production rates and operating costs, as well as environmental data from air quality sensors and greenhouse gas analyzers. The system demonstrates a 97.5% accuracy in evaluating corporate carbon performance, outperforming other models like K-Nearest Neighbors (KNN), Random Forest (RF), Decision Trees (DT), Naive Bayes (NB), and Logistic Regression (LR) [113].
@@
Zhao et al. proposed an innovative approach for predicting carbon emissions at the enterprise level, specifically targeting coal combustion processes in power generation organisations. This model utilizes Rough Set Theory to preprocess and extract key features from incomplete and imprecise carbon emission data, thereby enhancing the reliability of the data for subsequent modeling. The model then integrates a Grey System GM(1,1) model with a Support Vector Machine (SVM) to predict carbon emissions. The Grey model generates initial predictions based on historical data, which are further refined by the SVM to account for nonlinear relationships and improve accuracy. The input data for the model includes coal consumption rates, corresponding carbon emissions, and other related variables from a large coal-fired power company, covering multiple plants. The combination of the Grey model and SVM results in a hybrid prediction framework that significantly reduces prediction errors compared to traditional methods [114].
@@
Shi et al. developed a carbon emission prediction model tailored for thermal power enterprises, utilizing data from 17 power plants in Gansu Province, China. The study introduced a hybrid approach that combines Factor Analysis (FA) with multiple machine learning models, including Support Vector Regression (SVR) and a Genetic Algorithm-optimized Backpropagation Neural Network (GA-BP). The model inputs include critical variables such as energy consumption, electricity output, fuel type, and carbon content per unit of heat, derived from the 2021 greenhouse gas emissions reports and carbon verification documents of the power plants. Factor Analysis was employed to reduce data dimensionality by extracting key factors—such as energy consumption and output, energy quality, and energy efficiency—which were then used as inputs for the prediction models. Among the models evaluated, the GA-BP neural network demonstrated superior performance in predicting CO2 emissions, showcasing lower error rates and higher accuracy [115].
@@
Pan et al. introduced a real-time estimation method for the cement industry, utilizing a CNN-BLSTM (Convolutional Neural Network-Bidirectional Long Short-Term Memory) model. This method accurately estimates both direct and indirect carbon emissions by integrating CNN for feature extraction from material and energy consumption data, and BLSTM [116] for capturing temporal dependencies. Additionally, the study employs a Marginal Carbon Emission Factor (MCEF) to assess indirect emissions. The model was trained using 30 days of operational data from a cement corporation, demonstrating superior accuracy compared to traditional methods [117].
@@
After a thorough review of corporate carbon emission prediction algorithms, several conclusions can be drawn. The research in this field is increasingly moving towards more diverse data inputs, refined model processing, and industry-specific customization. By incorporating complex feature extraction techniques and multidimensional data inputs, researchers have enabled prediction models to more accurately capture the various factors influencing corporate carbon emissions. Moreover, the widespread application of machine learning and deep learning algorithms, along with the exploration of hybrid models and optimization algorithms, has significantly enhanced the accuracy and robustness of these prediction models. As the demand for real-time predictions grows, especially in high-emission industries, researchers have developed customized prediction models tailored to specific sectors, improving the effectiveness of these models, and Fig. 3 reflects the trends in the evolution of these technologies. These advancements not only assist organisations in better managing their carbon emissions but also provide investors and governments with more transparent and reliable carbon emission information, thereby supporting global efforts to achieve climate goals.

### v2

@@ 5 Algorithms for Predicting Carbon Credit Prices
[carbon38] LS-SVM Improved Particle Swarm Optimization (IPSO) EMD Y
[carbon39] LSTM, GRU, MLP, BPNN Genetic Algorithm (GA) CEEMDAN+Permutation Entropy+Variational Mode Decomposition N
[carbon40] KELM Multi-Objective Chaotic Sine Cosine Algorithm (MOCSCA) ICEEMDAN Y
[carbon41] SVR (high-frequency subsequences) + MLP (low-frequency subsequences) - ICEEMDAN+Discrete Wavelet Transform Y
[carbon42] BIGRU point prediction + GPR interval prediction - CEEMDAN+Wavelet Transform Y
[carbon43] GRU - Empirical Wavelet Transform N
[carbon44] Extreme Learning Machine, ELM Cooperative Kidney Algorithm Mean value Optimization Empirical Mode Decomposition, MOEMD Y
[carbon45] LSTM - MEEMD Y
[carbon47] XGBoost, RF, KNEA, SVM, RBFNN GWO, PSO, SA, FFOA CEEMDAN N
[carbon48] Weighted Regularized Extreme Learning Machine, WRELM Multi-Objective Grasshopper Optimization Algorithm, MOGOA ICEEMDAN Y
[carbon49] LSTM IWOA MEEMD N
[carbon51] VGM (long-term trend), ELM (short-term trend) PSO Hodrick-Prescott, HP filter Y
[np17] LSTM, Lower Upper Bound Estimation (LUBE) Asymmetric Multi-Objective Evolutionary Algorithm, MOEA CEEMDAN Y
[np16] Light Gradient Boosting Machine (LightGBM) Bayesian Optimization Algorithm, BOA CEEMDAN, Two-Stage Feature Selection, TFS Y
[np15] LSTM - Convolutional Neural Network N
[np14] LSTM - Multivariate Variational Mode Decomposition, MVMD Y
[np13] LSTM, GRU Cuckoo Search Algorithm, CS CEEMDAN, Fuzzy Entropy, FE, Variational Mode Decomposition Y
[np2] Deep Augmented Frequency Enhanced Decomposed Transformer (DA-FEDformer), Multimodel Optimization Piecewise Error Correction System - ICEEMDAN N
[xu2024novel] RELM (global and local sequence prediction), GPR (interval prediction) - Real-Time EMD, FIG N
[zheng2024multifactor] CatBoost (low-frequency), KELM (high-frequency) Singular Spectrum Analysis (SSA) Sequential Variational Mode Decomposition , Sample Entropy Y
[mao2024hybrid] LightGBM Grey Wolf Optimizer CEEMDAN, VMD, Sample Entropy Y
[mtemgnn2025] Multivariate Temporal Graph Neural Network (MtemGNN) - Dynamic Topic Model, sentiment feature construction Y
[qrtransformer_midas2025] Transformer-based quantile regression - Mixed-Data Sampling (MIDAS) Y
[cwtgpr_twicedec_2025] GRU Crested Porcupine Optimization (CPO) Complex Wavelet Transform, Gaussian Process twice decomposition N
Summarizing the research progress in carbon credit price prediction over the past few years, we can observe a gradual evolution and enhancement of the techniques employed. Early studies primarily relied on single models, such as the Hybrid Neural Fuzzy Controller (PATSOS), Adaptive Neuro-Fuzzy Inference System (ANFIS), and Artificial Neural Networks (ANN). These models played a crucial role in the initial stages of carbon price prediction [carbon46]. However, as data complexity increased, researchers began introducing data decomposition techniques, such as Empirical Mode Decomposition and its improved versions. These techniques allow time series data to be decomposed into more regular components, making it easier for machine learning algorithms to capture underlying patterns in the data. Additionally, optimization algorithms have been widely applied to enhance the overall performance of the models.
@@
To further improve prediction accuracy, recent research has concentrated on more refined data decomposition and optimized multi-model combinations. On the one hand, following data decomposition, researchers have employed more precise secondary decomposition techniques to extract potential patterns, enabling the models to learn and predict more accurately. On the other hand, the combination of different prediction models and optimization algorithms has been extensively explored to identify an effective hybrid model for different carbon markets. Moreover, some studies have decomposed raw data into components with different frequencies and applied different prediction algorithms to each frequency component, ultimately integrating the predictions of each sub-model through weighted ensemble methods. Meanwhile, some studies have introduced quantitative tools to measure the contribution of different factors to carbon price predictions, thereby enhancing the interpretability of the models.
In addition to point predictions, an increasing number of studies have introduced interval prediction techniques to provide a prediction range for carbon prices. This approach not only improves the reliability of predictions but also better addresses market uncertainties, offering more comprehensive predictive insights. Moreover, multi-step prediction techniques have gained increasing attention. This method recursively predicts carbon prices for multiple future time steps, capturing both short-term fluctuations and long-term trends. It provides more stable and comprehensive predictions when dealing with high market volatility and data uncertainties, thereby further enhancing the reliability and accuracy of predictions.
@@ 6 Prediction of Enterprise Carbon Emissions
Carbon emission prediction algorithms hold significant importance in the current global context of combating climate change, covering the broad range of carbons such as corporate operational carbon and building embodied carbon. Beyond assisting organizations in better managing their carbon emissions, these predictive methods are also crucial in meeting the needs of investors and governments. Financial institutions are increasingly focused on whether organizations adhere to ESG principles and prefer to invest in those that can create long-term societal value and impacts. Accurate carbon emission predictions help enhance corporate transparency, reduce information asymmetry, and strengthen market trust. Moreover, as governments impose stricter requirements on corporate carbon emission disclosures, precise carbon emission predictions can assist organizations in better complying with regulations, supporting government efforts to achieve emission reduction targets, and effectively evaluating whether organizations are truthfully disclosing and adhering to carbon emission rules. The following section will start with a review of various carbon emission prediction methods, ranging from the methods of empirical models and machine learning methods.
6.1 Empirical Models
@@
In the context of predicting corporate-level carbon emissions, the study by Goldhammer et al. presents an innovative approach that estimates corporate carbon footprints (CCFs) using externally available data. They employ a regression analysis model that primarily utilizes five key variables as input data: firm size (measured by turnover or number of employees), vertical integration, capital intensity, production centrality, and the carbon intensity of the national energy mix. The model was trained on actual CCF data from 93 European organisations across the chemical, construction engineering, and industrial machinery sectors. The study demonstrates that this method can accurately estimate corporate carbon emissions without relying on internal data, making it particularly suitable for scenarios requiring rapid, external assessments [carbon62].
@@
[carbon64] explores the application of two primary technical methods. The first is based on the Economic Input-Output Life Cycle Assessment (EIO-LCA) model, which utilizes macroeconomic data, such as industry GDP and carbon emission intensity, to estimate corporate carbon emissions. The second method is the estimation model developed by MSCI ESG Research, which combines historical company disclosures with industry-specific intensity to predict carbon emissions for years without disclosed data. The research used the 2013 carbon emission data from 277 organisations in the MSCI ACWI IMI index as training data and evaluated the models’ prediction accuracy by comparing them with the organisations’ actual disclosed data. The results indicate that the MSCI ESG Research model outperforms the EIO-LCA model in estimation accuracy, particularly in handling company-specific carbon emission intensities [carbon64].
@@
The study by Griffin et al. provides an effective technical approach for estimating corporate greenhouse gas (GHG) emissions using external data. This research employs the Ohlson valuation model [ohlson65] and combines it with reported GHG emission data to develop a GHG emission estimation model through regression analysis. This model utilizes a range of company-specific variables as input data, including revenue, capital expenditures, and asset composition, to predict the emissions of organisations that have not disclosed their GHG data [carbon63].
6.2 Machine Learning Methods
@@
Zhou et al. developed a carbon risk prediction model based on Support Vector Machines to accurately predict carbon emissions for heavily polluting industrial enterprises in China. The study first established a carbon risk assessment system comprising 45 initial indicators. After performing principal component analysis and independent analysis, 18 core indicators were selected. The SVM algorithm was used for model training, and the penalty parameter C and Gaussian kernel parameter gamma were optimized through grid search and cross-validation to handle the nonlinear and high-dimensional characteristics of the data [ep2].
@@
Quyen et al. developed a novel approach for predicting corporate-level carbon emissions, specifically targeting organisations that do not publicly disclose their emissions data. The framework proposed a Meta-Elastic Net learning model to aggregate predictions from multiple base models, including Ordinary Least Squares (OLS), Elastic Net, Neural Networks, K-Nearest Neighbors (KNN), Random Forest, and Extreme Gradient Boosting. The input training data comprised global carbon emissions data from 2,289 organisations between 2005 and 2017, totaling 13,435 annual observations. The data included Scope 1, Scope 2, and Scope 3 emissions, as well as various corporate factors such as operational scale, business models, technological advancements, and energy consumption. The model’s performance was validated through extensive cross-validation and robustness tests, demonstrating significant improvements in prediction accuracy compared to traditional regression models [carbon65].
@@
Li et al. proposed an advanced carbon emission prediction model tailored for enterprise-level applications, integrating multi-dimensional feature extraction with an optimized Radial Basis Function (RBF) neural network. The model utilizes Principal Component Analysis [pca] combined with kernel functions to extract essential features from a diverse set of input data, which includes production metrics, fuel consumption, electricity and heat energy purchases, and other relevant economic indicators. This approach effectively reduces data complexity while preserving key predictive information. The extracted features are then input into an RBF neural network, whose parameters are finely tuned using a hybrid optimization method that combines the Drosophila Optimization Algorithm and Differential Evolution. This dual optimization strategy significantly enhances the accuracy and robustness of the emission predictions [ep3].
@@
Chen et al. proposed an efficient Fast Transfer-Based Adaptive Recurrent Neural Network (SEER) model aimed at addressing the high computational costs associated with real-time predictions in existing deep learning models. The SEER model enhances prediction accuracy and training efficiency by integrating data preprocessing, feature extraction, attention mechanisms, transfer learning, and adaptive learning rate mechanisms. Using carbon emission and air quality data from Shanghai and Beijing as inputs, the model effectively identifies key data points and optimizes generalization capabilities, demonstrating its accuracy and applicability in corporate carbon emission predictions [ep7].
@@
Tang et al. proposed CarbonNet, a novel approach for enterprise-level carbon emission prediction, aimed at overcoming the limitations of traditional macro-level forecasting models. The paper first constructs a large-scale dataset encompassing 3346 organisations over 31 years, integrating carbon emission data with corporate financial and operational metrics. To enhance prediction accuracy, the authors employed Factor Analysis for feature extraction, reducing the impact of irrelevant variables, thereby building a more generalizable carbon emission prediction model. Subsequently, the paper utilized the XGBoost algorithm for large-scale data mining, iteratively optimizing prediction errors [ep1].
@@
Shou developed a corporate carbon performance evaluation framework using Support Vector Machine to predict carbon emissions at the enterprise level. The framework, named Enterprise Carbon Performance Index System (ECPIS), integrates SVM with real-time data collection and monitoring to assess and forecast carbon emissions, enabling organisations to make data-driven decisions to reduce their carbon footprint. The input training data for the SVM model includes historical carbon emissions data, energy consumption metrics, operational data such as production rates and operating costs, as well as environmental data from air quality sensors and greenhouse gas analyzers. The system demonstrates a high accuracy in evaluating corporate carbon performance, outperforming other models like K-Nearest Neighbors, Random Forest, Decision Trees (DT), Naive Bayes (NB), and Logistic Regression [ep5].
@@
Zhao et al. proposed an innovative approach for predicting carbon emissions at the enterprise level, specifically targeting coal combustion processes in power generation organisations. This model utilizes Rough Set Theory to preprocess and extract key features from incomplete and imprecise carbon emission data, thereby enhancing the reliability of the data for subsequent modeling. The model then integrates a Grey System GM(1,1) model with a Support Vector Machine to predict carbon emissions. The Grey model generates initial predictions based on historical data, which are further refined by the SVM to account for nonlinear relationships and improve accuracy. The input data for the model includes coal consumption rates, corresponding carbon emissions, and other related variables from a large coal-fired power company, covering multiple plants. The combination of the Grey model and SVM results in a hybrid prediction framework that significantly reduces prediction errors compared to traditional methods [ep4].
@@
Shi et al. developed a carbon emission prediction model tailored for thermal power enterprises, utilizing data from 17 power plants in Gansu Province, China. The study introduced a hybrid approach that combines Factor Analysis (FA) with multiple machine learning models, including Support Vector Regression and a Genetic Algorithm-optimized Backpropagation Neural Network (GA-BP). The model inputs include critical variables such as energy consumption, electricity output, fuel type, and carbon content per unit of heat, derived from the 2021 greenhouse gas emissions reports and carbon verification documents of the power plants. Factor Analysis was employed to reduce data dimensionality by extracting key factors—such as energy consumption and output, energy quality, and energy efficiency—which were then used as inputs for the prediction models. Among the models evaluated, the GA-BP neural network demonstrated superior performance in predicting CO2 emissions, showcasing lower error rates and higher accuracy [ep6].
@@
Pan et al. introduced a real-time estimation method for the cement industry, utilizing a CNN-BLSTM (Convolutional Neural Network–Bidirectional Long Short-Term Memory) model. This method accurately estimates both direct and indirect carbon emissions by integrating CNN for feature extraction from material and energy consumption data, and BLSTM [Blstmep8] for capturing temporal dependencies. Additionally, the study employs a Marginal Carbon Emission Factor (MCEF) to assess indirect emissions. The model was trained using 30 days of operational data from a cement corporation, demonstrating superior accuracy compared to traditional methods [ep8].
@@
After a thorough review of corporate carbon emission prediction algorithms, several conclusions could be drawn. The research in this field is increasingly moving towards more diverse data inputs, refined model processing, and industry-specific customization. By incorporating complex feature extraction techniques and multidimensional data inputs, researchers have enabled prediction models to more accurately capture the various factors influencing corporate carbon emissions. Moreover, the widespread application of machine learning and deep learning algorithms, along with the exploration of hybrid models and optimization algorithms, has significantly enhanced the accuracy and robustness of these prediction models. As the demand for real-time predictions grows, especially in high-emission industries, researchers have developed customized prediction models tailored to specific sectors, improving the effectiveness of these models, and Fig. 3 reflects the trends in the evolution of these technologies. These advancements not only assist organisations in better managing their carbon emissions but also provide investors and governments with more transparent and reliable carbon emission information, thereby supporting global efforts to achieve climate goals.
6.3 Predictive Models for Embodied Carbon Emission
In addition to carbon emissions generated during corporate operations, carbon emissions associated with buildings owned or utilized by enterprises should also be considered. Data indicates that emissions from building-related activities account for more than 39% of global emissions, with embodied carbon contributing approximately 11% [embodiedcarbon2]. Therefore, accurately quantifying the embodied carbon within buildings is essential for comprehensively predicting carbon emissions and correctly attributing them to responsible enterprises.
Zhang et al. collected data from 850 building samples across China, covering structural characteristics and material usage. The dataset included 13 structural features such as building height, structural type, seismic design intensity, geographic location, delivery method, project cost, and material expenses. Additionally, 16 material-related features were recorded, including the consumption of steel, concrete, bricks, prefabricated components, doors, windows, and paints. The study applied a process-based Life Cycle Assessment methodology to calculate the Embodied Carbon Intensity during the material production stage. Outliers were identified and removed using the Isolation Forest algorithm, followed by feature selection through Spearman’s correlation coefficient analysis to eliminate highly correlated variables. The researchers then employed 12 machine learning models for prediction, categorized as follows: linear models, Ridge Regression (RD), Bayesian Ridge Regression (BR), Elastic Net Regression (EN); single nonlinear models, K-Nearest Neighbors, Support Vector Regression; and ensemble models like Gradient Boosting, Extreme Gradient Boosting (XGB), Histogram-based Gradient Boosting, Random Forest, and Extremely Randomized Trees. The ET model demonstrated superior performance during the planning and design phase, with building height, material cost, and geographic region identified as the core predictive features. During the preliminary design stage, the XGB model achieved the highest performance, with material costs, rebar, and concrete usage emerging as the most influential factors [embodiedcarbon1].
Su et al. utilized data from 70 construction projects in the Yangtze River Delta region of China, capturing various building and construction-related characteristics. The data underwent preprocessing using the Quartile Method to eliminate outliers, followed by KNN-based imputation for missing values. Spearman’s correlation analysis was employed to remove redundant features, ultimately yielding 15 project-level attributes (e.g., building height, location, structural type) and 12 construction-level attributes (e.g., on-site workforce, prefabricated concrete volume). SVR, ANN, and XGBoost models were applied to predict material consumption, which was then multiplied by corresponding Emission Factors to estimate total embodied carbon emissions. The model achieved a prediction error of less than 5.33% [embodiedcarbon3].
Marsh et al. introduced a Monte Carlo Simulation–based approach to quantify uncertainties in embodied carbon (EC) calculations during the building product stage. The methodology begins by extracting material consumption data from construction inventories and carbon emission coefficients from established databases such as ICE and OneClickLCA. Outliers were removed using the Quartile Method, while Spearman’s correlation analysis helped exclude redundant features. The study assumed that Embodied Carbon Coefficients (ECCs) follow a normal distribution, with missing values imputed based on the global mean coefficient of variation (CoV). Subsequently, 50,000 Monte Carlo simulations were conducted to generate a probabilistic distribution of EC predictions across different scenarios. The results included mean values, standard deviations, and confidence intervals to evaluate error margins resulting from data uncertainty. Sensitivity analysis revealed that high-variance materials, particularly rebar and steel, substantially influenced the prediction’s dispersion. The findings emphasized that neglecting ECC uncertainty could lead to prediction deviations exceeding 10%, underscoring the importance of incorporating uncertainty analysis during early design stages to enhance prediction accuracy and decision-making reliability [embodiedcarbon4].
For a comprehensive prediction of corporate carbon emissions, embodied carbon also plays a crucial role. Focusing solely on emissions generated during corporate operations while neglecting the carbon emissions embedded in buildings and infrastructure can lead to a systematic underestimation of the company’s overall carbon footprint, ultimately compromising the scientific rigor and effectiveness of carbon management and policy-making. Currently, the primary method for predicting embodied carbon emissions in buildings is the Carbon Emission Factor approach, which involves identifying the individual components of building materials and assigning a corresponding carbon emission factor to each, thereby quantifying the carbon footprint of the building throughout its construction and life cycle stages. However, these carbon emission factors are influenced by various dynamic factors, such as production processes, material sources, and supply chain fluctuations, resulting in temporal variability. Consequently, researchers have increasingly adopted carbon emission uncertainty quantification methods, such as Monte Carlo simulations, to assess the potential prediction deviations, establishing this as a key research direction in the field. By systematically analyzing and predicting the embodied carbon characteristics in buildings, these research advancements not only support enterprises in formulating scientifically sound carbon management strategies but also provide critical insights that help market participants and regulatory bodies make informed decisions, ultimately contributing to the collective global response to climate change.